solve the system of equations
4x-5y=-14
x=3y

Found 2 solutions by flame8855, lynnlo:
Answer by flame8855(424)

4(3y) - 5y = -14
13y-5y = -14
y= 7/4
x = 21/4

Answer by lynnlo(4176)

y=6√3402/486

y=-6√3402/486